OpenStudy (anonymous):

110 Kg/in^2 to oz/um^2

ganeshie8 (ganeshie8):

they mixing english units with metric units

OpenStudy (anonymous):

(110 kg) per (square inches) = 6.01422254 × 10-6 ounces per (square micrometer)

OpenStudy (anonymous):

umm um=micrometers

OpenStudy (anonymous):

google

OpenStudy (anonymous):

@josiahh can you make that less confusing

OpenStudy (anonymous):

I googled it and that is the answer it gave me

OpenStudy (anonymous):

can you send me the link?

OpenStudy (anonymous):

just copy and paste this to google: 110 kg per square inches to ounces per square micrometer

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

@josiahh can you help with 3600.0 m/s^2= Kmi/day^2

OpenStudy (anonymous):

what's Kmi?

OpenStudy (anonymous):

kilo and miles

OpenStudy (anonymous):

1km=0.62137119

OpenStudy (anonymous):

miles

OpenStudy (anonymous):

ok so 1 kilomile is 1000 miles...

OpenStudy (anonymous):

right...

ganeshie8 (ganeshie8):

google wont help this time, we need to get familiar wid dimensional analysis, it would help u later in your course when u start doing more complex problems

OpenStudy (anonymous):

divide the answer from this link by 1000 and you have your kmi/days^2

ganeshie8 (ganeshie8):

!

OpenStudy (anonymous):

it should look something like this